# Interference Among First-Price Pacing Equilibria: A Bias and Variance Analysis

### v1

@@ 2.2.1. Basic Definitions, Limit FPPE and Regularity Conditions
Let $\beta^{*}$ and $p^{*}$ be the equilibrium pacing multipliers and prices. Revenue in the limit FPPE is ${\mathsf{{REV}}}^{*}=\int p^{*}(\theta)s(\theta)\mathop{}\!\mathrm{d}\theta\;.$ It measures the profitability of the auction platform. The leftover budgets for buyers are ${\delta^{*}_{i}}=b_{i}-\int p^{*}(\theta)s(\theta){x^{*}_{i}}(\theta)\mathop{}\!\mathrm{d}\theta$.
The first two conditions simply describe the possible outcomes of an auction system that uses pacing as the budget management strategy. The last condition, no unnecessary pacing, ensures that we only scale down a buyer’s bids in case their budget constraint is binding. FPPE has many nice properties, including that they are competitive equilibria and that they are revenue-maximizing among budget-feasible pacing multipliers (Conitzer et al., 2022a).
In a pacing auction market ${\mathscr{M}}={\small\mathsf{{FPPE}}}(b,v,s,\Theta)$ the following two regularity conditions are important for the study of its statistical properties.
@@
We say the smoothness condition holds if the map $\beta\mapsto{\mathbb{E}}_{s}[\max_{i}{\beta_{i}}{v_{i}(\theta)}]$ is twice continuously differentiable in a neighborhood of $\beta^{*}$.
@@
The condition Definition 2 (SMO). ensures that in the limit market, items that incur a tie are measure zero. The condition Definition 3 (SCS). rules out degenerate buyers that spend their budget exactly at $\beta_{i}^{*}$. See Liao and Kroer (2023) and Liao et al. (2023) for an extensive discussion about these conditions.
2.2.2. Finite FPPE
Next we introduce finite FPPE, which models the auction data we observe in practice. Let $\gamma=(\theta^{1},\dots,\theta^{t})$ be a sequence of items. Assume each item has the same supply of $\sigma\in{\mathbb{R}_{+}}$ units. A finite FPPE is a limit FPPE where the supply is a discrete measure supported on the observed items $\gamma$. Let ${v_{i}^{\tau}}={v_{i}(\theta^{\tau})}$.
Definition 4 (Finite FPPE, Conitzer et al. (2022a)).
Given $(b,v,\sigma,\gamma)$, the observed FPPE, ${\small\widehat{\textsf{{FPPE}}}}(b,v,\sigma,\gamma)$, is the unique tuple $(\beta,p)\in[0,1]^{n}\times{\mathbb{R}}^{t}_{+}$ such that there exist $x_{i}^{\tau}\in[0,1]$:
  • (First-price) For all $\tau$, ${p^{\tau}}=\max_{i}{\beta_{i}}{v_{i}^{\tau}}$. For all $i$ and $\tau$, ${x_{i}^{\tau}}>0$ implies ${\beta_{i}}{v_{i}^{\tau}}=\max_{k}\beta_{k}v_{k}^{\tau}$.
  • (Supply and budget feasible) For all $i$, $\sigma{\sum_{\tau=1}^{t}}{x_{i}^{\tau}}{p^{\tau}}\leq b_{i}$. For all $\tau$, ${\sum_{i=1}^{n}}{x_{i}^{\tau}}\leq 1$. (Market clearing) For all $\tau$, ${p^{\tau}}>0$ implies ${\sum_{i=1}^{n}}{x_{i}^{\tau}}=1$.
  • (No unnecessary pacing) For all $i$, $\sigma{\sum_{\tau=1}^{t}}{x_{i}^{\tau}}{p^{\tau}}<b_{i}$ implies ${\beta_{i}}=1$.
In Liao and Kroer (2023), it is shown that if $\gamma$ consists of $t$ i.i.d. draws from distribution $s$, and one takes $\sigma=1/t$, then the pacing multiplier in ${\small\widehat{\textsf{{FPPE}}}}(b,v,1/t,\gamma)$ converges to the pacing multiplier in ${\small\mathsf{{FPPE}}}(b,v,s,\Theta)$ in probability. Also note that the ${\small\widehat{\textsf{{FPPE}}}}(tb,v,1,\gamma)$ converges to the same limit FPPE ${\small\mathsf{{FPPE}}}(b,v,s,\Theta)$ because the pacing multiplier of a finite FPPE does not change when budgets and supplies are multiplied by the same scalar.
2.2.3. The Eisenberg-Gale Program
Both the limit FPPE and the finite FPPE have convex program characterizations (Chen et al., 2007; Conitzer et al., 2022a; Gao and Kroer, 2022). We define the dual Eisenberg-Gale (EG) objective for a single item $\theta$ as
(1) $F(\theta,\beta)=f(\beta,\theta)-{\sum_{i=1}^{n}}b_{i}\log\beta_{i}\;,\quad f(\beta,\theta)=\max_{i\in[n]}\beta_{i}v_{i}(\theta)\;.$
@@ 2.2.3. The Eisenberg-Gale Program
(2) $H(\beta)={\mathbb{E}}[F(\theta,\beta)]\;,\quad H_{t}(\beta)=\frac{1}{t}{\sum_{\tau=1}^{t}}F({\theta^{\tau}},\beta)\;.$
@@
We say ${\mathscr{H}}$ is the Hessian of market ${\mathscr{M}}$ when ${\mathscr{H}}=\nabla^{2}\int F(\theta,\beta)s\mathop{}\!\mathrm{d}\theta|_{\beta=\beta^{*}}$.
@@
The equilibrium pacing multipliers $\beta^{*}$ in ${\small\mathsf{{FPPE}}}(b,v,s,\Theta)$ can be recovered through the population dual EG program
(3) $\beta^{*}=\operatorname*{arg\,min}_{\beta\in(0,1]^{n}}H(\beta)\;.$
@@
The pacing multiplier $\beta^{*}$ is the unique solution to Eq. 3. Let $\beta^{\gamma}$ be the equilibrium pacing multiplier in ${\small\widehat{\textsf{{FPPE}}}}(b,v,1/t,\gamma)$. Then $\beta^{\gamma}$ solves the sample analogue of Eq. 3:
(4) $\beta^{\gamma}=\operatorname*{arg\,min}_{\beta\in(0,1]^{n}}H_{t}(\beta)\;.$
@@
Let us briefly comment on the differential structure of $f$, since it plays a role in later sections. The function $f(\beta,\theta)$ is a convex function of $\beta$ and its subdifferential $\partial_{\beta}f(\beta,\theta)$ is the convex hull of $\{v_{i}e_{i}:\text{index $i$ such that ${\beta_{i}}{v_{i}(\theta)}=\max_{k}{\beta_{k}}v_{k}(\theta)$}\}$, with $e_{i}$ being the base vector in ${\mathbb{R}^{n}}$. When $\max_{i}{\beta_{i}}v_{i}(\theta)$ is attained by a unique $i^{*}$, the function $f(\cdot,\theta)$ is differentiable. In that case, all entries of $\nabla_{\beta}f(\beta,\theta)$ are zero expect that the $i^{*}$-th entry is filled with the value $v_{i^{*}}(\theta)$.
3. Setup: Interference as Contamination
@@
In this section, we discuss how to estimate market equilibria when there is contamination in the supply. Then, we show that interference from other markets can be viewed as a form of contamination, and so the problem of removing interference bias can be analyzed via our contamination framework.
3.1. Contaminated First-Price Pacing Equilibrium
We assume that we are in the same FPPE setting as before: there are $n$ buyers, each with budget $b_{i}$, and an item set $\Theta$. However, now we assume that $\Theta$ is partitioned into $\Theta_{\mathsf{good}}$ and $\Theta_{\mathsf{bad}}$. Let $s$ be supported on $\Theta_{\mathsf{good}}$ and $s^{\prime}$ on $\Theta_{\mathsf{bad}}$. We define the $\alpha$-contaminated market and the uncontaminated market by
${\mathscr{M}}_{\alpha}={\small\mathsf{{FPPE}}}(b,v,s_{\alpha},\Theta)\;,\quad{\mathscr{M}}_{0}={\small\mathsf{{FPPE}}}(b,v,s,\Theta)\;,$
where $s_{\alpha}=\alpha s^{\prime}+(1-\alpha)s$. Our goal is to perform inference about FPPE properties in the limit FPPE with the item set $\Theta_{\mathsf{good}}$ and supply $s$. However, we are given access to finite FPPEs sampled from $s_{\alpha}$ instead. In particular, let $\gamma$ be $t$ i.i.d. draws from $s_{\alpha}$ and let
^$\widehat{\mathscr{M}}_{\alpha}={\small\widehat{\textsf{{FPPE}}}}(b,v,1/t,\gamma)\;.$
@@ 3.1. Contaminated First-Price Pacing Equilibrium
We assume knowledge of $\alpha$ throughout the paper. In practice, this can often be estimated from historical data; in the parallel A/B test setting, this can be estimated directly from the sampled set of items, since we know which items are good and bad in that setting. Let $\beta^{*}$ and $\beta^{*}_{\alpha}$ be the limit pacing multipliers in ${\mathscr{M}}_{0}$ and ${\mathscr{M}}_{\alpha}$, respectively. Let $\beta^{\gamma}_{\alpha}$ be the pacing multipliers in the sampled market ^$\widehat{\mathscr{M}}_{\alpha}$ and let $H_{\alpha,t}(\beta)=\frac{1}{t}{\sum_{\tau=1}^{t}}F({\theta^{\tau}},\beta)$ be the sample EG objective.
@@
Now, if we wanted to make inferences about ${\mathscr{M}}_{\alpha}$ then we could use existing statistical inference theory on how to use data in a finite FPPE ^$\widehat{\mathscr{M}}_{\alpha}$ to make inferences about the limit FPPE ${\mathscr{M}}_{\alpha}$. In recent work of Liao and Kroer (2023), it is shown that for an influence function $D_{\alpha}:\Theta\to{\mathbb{R}}$ (to be defined in Thm. 3), the asymptotic distribution of $\beta_{\alpha}^{\gamma}$ around $\beta_{\alpha}^{*}$ satisfies
$\sqrt{t}(\beta^{\gamma}_{\alpha}-\beta^{*}_{\alpha})=\frac{1}{\sqrt{t}}{\sum_{\tau=1}^{t}}(D_{\alpha}({\theta^{\tau}})-{\mathbb{E}}_{\alpha}[D_{\alpha}(\theta)])+o_{p}(1)\;,$
where $\{{\theta^{\tau}}\}$ is the set of items sampled in $\gamma$. Our central research question can now be formulated as follows:
@@
How can we use the data in the finite contaminated market ^$\widehat{\mathscr{M}}_{\alpha}$ to make inferences about the uncontaminated limit market ${\mathscr{M}}_{0}$?
3.2. Interference as Contamination
@@
Now we show how the contamination model from the previous section can be used to model interference. Consider $K$ separated auction markets, which together form a global market. In the global market there are $n$ buyers, each with budget $b_{i}$, and an item set $\Theta$, partitioned into $\Theta_{\mathsf{good}}$ and $\Theta_{\mathsf{bad}}$. Let $C_{1},\dots,C_{K}$ be a partition of the buyers, $\Theta_{1},\dots,\Theta_{K}$ be a partition of the good item set $\Theta_{\mathsf{good}}$, and $s_{1},\dots s_{K}$ be a set of supply functions, supported on $\Theta_{1},\dots,\Theta_{K}$ respectively. The $k$-th submarket consists of buyers in $C_{k}$, the item set $\Theta_{k}$ and supply $s_{k}$. Let $s=\frac{1}{K}\sum_{k}s_{k}$ be the average mixture and $s^{\prime}$ be a supply supported on $\Theta_{\mathsf{bad}}$. Let the contaminated supply be $s_{\alpha}=\alpha s^{\prime}+(1-\alpha)s$.
@@
In Fig 3 we present an example of interference among $K=3$ submarkets. The market of interest is the perfectly separated market (right). The reason is, in parallel A/B testing, submarkets are created in a way that each resembles the global market, and when a submarket receives a treatment, the observed quantities in that submarket, such as revenues and social welfare, are considered surrogates for the treatment effect in the global market. However, in practice we only observe the interfered finite market (left), which converges to the interfered limit market (middle). In Sec. 5 we show how to analyze parallel A/B testing using this framework.
Figure 3. The model. Left (^$\widehat{\mathscr{M}}_{\alpha}$): the observed market where interference is present among submarkets. Middle (${\mathscr{M}}_{\alpha}$): the limit market with interference from bad item set. Right (${\mathscr{M}}_{0}$): the limit market with perfectly separated submarkets. We use data from the left panel to make inferences about the market in the right panel.
4. Main Results
@@
In this section, we develop a methodology for making inferences about, e.g., pacing multipliers or revenue in the uncontaminated limit FPPE. We introduce a surrogate for pacing multipliers, based on the notion of directional derivatives, and establish its debiasing property in Sec. 4.1. Then, we focus on estimating this surrogate quantity in Sec. 4.2, and develop asymptotic normality results in Sec. 4.3. Secondly, we consider estimating revenue. Revenue can be thought of as a smooth function of pacing multipliers, and we discuss debiased estimation and inference of revenue based on our pacing multiplier results in Sec. 4.4.
Before we proceed we briefly discuss two baseline approaches to the problem; see Table 2 for an overview. The first method is doing nothing, i.e., treating quantities in the contaminated finite market ^$\widehat{\mathscr{M}}_{\alpha}$ as estimators of those in ${\mathscr{M}}_{0}$. This approach suffers from a significant level of bias, roughly of order $\Theta(\alpha)$, and the bias does not vanish as the number of items increases. The second method is that, since we know whether an item comes from the contamination supply $s^{\prime}$, we could throw away the market equilibrium ^$\widehat{\mathscr{M}}_{\alpha}$, collect the good items, and compute a finite market equilibrium formed by the good items via solving the EG program in Eq. 4. This approach has a vanishing bias but is computationally heavy, especially in real-world advertising platforms where there are millions of buyers. Our proposed approach reduces bias to the order $o(\alpha)$ and does not require solving large-scale EG programs, by utilizing the equilibrium structure of the observed market ^$\widehat{\mathscr{M}}_{\alpha}$.
Table 2. Comparison of Procedures
Procedure Bias Requires solving EG
Ignore interference $\Theta(\alpha)$ No
Solve EG $o(1)$ Yes
Our debiasing procedure $o(\alpha)$ No
4.1. A Debiased Surrogate for Pacing Multipliers

### v3

@@ 2 Review of FPPE Theory
Let ${\bm{\beta}}^{*}$ and $p^{*}$ be the equilibrium pacing multipliers and prices. Revenue in the limit FPPE is ${{{REV}}}^{*}=\int p^{*}(\theta)s(\theta)\mathop{}\!\mathrm{d}\theta\;.$ It measures the profitability of the auction platform. The leftover budgets for buyers are denoted by ${\delta^{*}_{i}}=b_{i}-\int p^{*}(\theta)s(\theta){x^{*}_{i}}(\theta)\mathop{}\!\mathrm{d}\theta$.
The first two conditions simply describe the possible outcomes of a first-price auction system that uses pacing as the budget management strategy. The last condition, no unnecessary pacing, ensures that we only scale down a buyer’s bids in case their budget constraint is binding. FPPE has many nice properties, including that they are competitive equilibria and that they are revenue-maximizing among budget-feasible pacing multipliers (Conitzer et al., 2022a).
In a pacing auction market ${\mathscr{M}}={\mathsf{{FPPE}}}(\bm{b},\bm{v},s,\Theta)$ the following two regularity conditions are important for the study of its statistical properties.
@@
We say the smoothness condition holds if the map ${\bm{\beta}}\mapsto{\mathbb{E}}_{s}[\max_{i}{\beta_{i}}{v_{i}(\theta)}]$ is twice continuously differentiable in a neighborhood of ${\bm{\beta}}^{*}$.
@@
The condition (SMO). ensures that in the limit market, items that incur a tie are measure zero. The condition (SCS). rules out degenerate buyers that spend their budget exactly at $\beta_{i}^{*}$. See Liao & Kroer (2023) and Liao et al. (2023) for an extensive discussion about these conditions.
Finite FPPE. Next we introduce the finite FPPE, which models the auction data we observe in practice. Let $\gamma=(\theta^{1},\dots,\theta^{t})$ be a sequence of items. Assume each item has the same supply of $\sigma\in{\mathbb{R}_{+}}$ units. A finite FPPE is a limit FPPE where the supply is a discrete measure supported on the observed items $\gamma$. Let ${v_{i}^{\tau}}={v_{i}(\theta^{\tau})}$.
Definition 4 (Finite FPPE, informal).
A finite FPPE, ${\widehat{\textsf{{FPPE}}}}(\bm{b},\bm{v},\sigma,\gamma)$, is a limit FPPE where the item set is the finite set of observed items $\gamma$. See App. C for the full definition.
In Liao & Kroer (2023), it is shown that if $\gamma$ consists of $t$ i.i.d. draws from distribution $s$, and one takes $\sigma=1/t$, then the pacing multiplier in ${\widehat{\textsf{{FPPE}}}}(\bm{b},\bm{v},1/t,\gamma)$ converge to the pacing multiplier in ${\mathsf{{FPPE}}}(\bm{b},\bm{v},s,\Theta)$ in probability. Also, note that the FPPE ${\widehat{\textsf{{FPPE}}}}(t\bm{b},\bm{v},1,\gamma)$ converges to the same limit FPPE ${\mathsf{{FPPE}}}(\bm{b},\bm{v},s,\Theta)$ because the pacing multipliers of a finite FPPE do not change when budgets and supplies are multiplied by the same scalar.
The Eisenberg-Gale Program. Both the limit FPPE and the finite FPPE have convex program characterizations (Chen et al., 2007; Conitzer et al., 2022a; Gao & Kroer, 2022). We define the dual Eisenberg-Gale (EG) objective for a single item $\theta$ as
$F(\theta,{\bm{\bm{\beta}}})=f(\theta,{\bm{\bm{\beta}}})-{\sum_{i=1}^{n}}b_{i}\log\beta_{i}\;,\quad f(\theta,{\bm{\bm{\beta}}})=\max_{i\in[n]}\beta_{i}v_{i}(\theta)\;.$ (1)
@@
$H({\bm{\beta}})={\mathbb{E}}[F(\theta,{\bm{\bm{\beta}}})]\;,\quad H_{t}({\bm{\beta}})=\frac{1}{t}{\sum_{\tau=1}^{t}}F({\theta^{\tau}},{\bm{\bm{\beta}}})\;.$ (2)
@@
We say ${\bm{H}}$ is the Hessian of market ${\mathscr{M}}$ when ${\bm{H}}=\nabla_{{\bm{\bm{\beta}}}{\bm{\bm{\beta}}}}^{2}\int F(\theta,{\bm{\beta}})s\mathop{}\!\mathrm{d}\theta|_{{\bm{\beta}}={\bm{\beta}}^{*}}$.
@@
The equilibrium pacing multipliers ${\bm{\beta}}^{*}$ in ${\mathsf{{FPPE}}}(\bm{b},\bm{v},s,\Theta)$ can be recovered through the population dual EG program
${\bm{\beta}}^{*}=\operatorname*{arg\,min}_{{\bm{\beta}}\in(0,1]^{n}}H({\bm{\beta}})\;.$ (3)
@@
The pacing multiplier vector ${\bm{\beta}}^{*}$ is the unique solution to Eq. 3. Let ${\bm{\beta}}^{\gamma}$ be the equilibrium pacing multiplier in ${\widehat{\textsf{{FPPE}}}}(\bm{b},\bm{v},1/t,\gamma)$. Then ${\bm{\beta}}^{\gamma}$ solves the sample analogue of Eq. 3:
${\bm{\beta}}^{\gamma}=\operatorname*{arg\,min}_{{\bm{\bm{\beta}}}\in(0,1]^{n}}H_{t}({\bm{\bm{\beta}}})\;.$ (4)
@@
Let us briefly comment on the differential structure of $f$, since it plays a role in later sections. The function $f({\bm{\bm{\beta}}},\theta)$ is a convex function of ${\bm{\bm{\beta}}}$ and its subdifferential $\partial_{\bm{\bm{\beta}}}f({\bm{\bm{\beta}}},\theta)$ is the convex hull of $\{v_{i}\bm{e}_{i}:\text{index $i$ such that ${\beta_{i}}{v_{i}(\theta)}=\max_{k}{\beta_{k}}v_{k}(\theta)$}\}$, with $\bm{e}_{i}$ being the base vector in ${\mathbb{R}^{n}}$. When $\max_{i}{\beta_{i}}v_{i}(\theta)$ is attained by a unique $i^{*}$, the function $f(\cdot,\theta)$ is differentiable. In that case, all entries of $\nabla_{\bm{\bm{\beta}}}f({\bm{\bm{\beta}}},\theta)$ are zero expect that the $i^{*}$-th entry is filled with the value $v_{i^{*}}(\theta)$.
3 Interference as Contamination
@@
In this section, we discuss how to estimate market equilibria when there is contamination in the supply, meaning that items are generated from a mixture of two distributions, when in reality we wish to estimate equilibrium quantities from one of the two distributions. Then, we show that interference from other markets can be viewed as a form of contamination, and so the problem of removing interference bias can be analyzed via our contamination framework.
3.1 FPPE with Contaminated Supply
@@
We assume that we are in the same FPPE setting as before: there are $n$ buyers, each with budget $b_{i}$, and an item set $\Theta$ which is now partitioned in to $\Theta_{\mathsf{bad}}$ and $\Theta_{\mathsf{good}}$. However, now we assume that the supply $s$ is contaminated by $s^{\prime}$, another supply distribution. We define the $\alpha$-contaminated market as ${\mathscr{M}}_{\alpha}={\mathsf{{FPPE}}}(\bm{b},\bm{v},s_{\alpha},\Theta)$ and the uncontaminated market as ${\mathscr{M}}_{0}={\mathsf{{FPPE}}}(\bm{b},\bm{v},s,\Theta)$, where $s_{\alpha}=\alpha s^{\prime}+(1-\alpha)s$, distribution $s^{\prime}$ is supported on $\Theta_{\mathsf{bad}}$, and $s$ on $\Theta_{\mathsf{good}}$. Our goal is to perform inference about FPPE properties in the limit FPPE with the supply $s$. However, we are given access to finite FPPEs sampled from $s_{\alpha}$ instead. In particular, let $\gamma$ be $t$ i.i.d. draws from $s_{\alpha}$ and let ^$\widehat{\mathscr{M}}_{\alpha}={\widehat{\textsf{{FPPE}}}}(\bm{b},\bm{v},1/t,\gamma)$. We assume $\alpha$ is known throughout the paper. In practice, this can often be estimated from historical data; in the parallel A/B test setting, this can be estimated directly from the sampled set of items, since we know whether an item is drawn from $s$ or $s^{\prime}$. Let ${\bm{\beta}}^{*}$ and ${\bm{\beta}}^{*}_{\alpha}$ be the limit pacing multipliers in ${\mathscr{M}}_{0}$ and ${\mathscr{M}}_{\alpha}$, respectively. Let ${\bm{\beta}}^{\gamma}_{\alpha}$ be the pacing multipliers in the sampled market ^$\widehat{\mathscr{M}}_{\alpha}$ and let $H_{\alpha,t}({\bm{\bm{\beta}}})=\frac{1}{t}{\sum_{\tau=1}^{t}}F({\theta^{\tau}},{\bm{\bm{\beta}}})$ be the sample EG objective.
@@
If we wanted to make inferences about ${\mathscr{M}}_{\alpha}$ then we could use existing statistical inference theory on how to use data in a finite FPPE ^$\widehat{\mathscr{M}}_{\alpha}$ to make inferences about the limit FPPE ${\mathscr{M}}_{\alpha}$ (Liao & Kroer, 2023). However, the supply contamination prevents the application of these tools to our problem.
@@
Our central research question is then on how to use data in the finite contaminated market ^$\widehat{\mathscr{M}}_{\alpha}$ to make inferences about the uncontaminated limit market ${\mathscr{M}}_{0}$.
In Sec. 4 we propose an estimator for this problem and derive its properties. The results there apply to general item space $\Theta$ and supplies $s$ and $s^{\prime}$. By imposing structure on $\Theta,s$ and $s^{\prime}$, we show that the contamination model captures the interference among FPPEs.
3.2 Application: Modeling Interference among FPPEs
@@
Now we show how the contamination model from the previous section can be used to model interference. Consider $K$ separate auction markets, which together form a global market. In the global market there are $n$ buyers, each with budget $b_{i}$, and an item set $\Theta$, partitioned into $\Theta_{\mathsf{good}}$ and $\Theta_{\mathsf{bad}}$. Let $C_{1},\dots,C_{K}$ be a partition of the buyers, $\Theta_{1},\dots,\Theta_{K}$ be a partition of the good item set $\Theta_{\mathsf{good}}$, and $s_{1},\dots s_{K}$ be a set of supply functions, supported on $\Theta_{1},\dots,\Theta_{K}$ respectively. The $k$-th submarket consists of buyers in $C_{k}$, the item set $\Theta_{k}$ and supply $s_{k}$. Let $s=\frac{1}{K}\sum_{k}s_{k}$ be the average mixture and $s^{\prime}$ be a supply supported on $\Theta_{\mathsf{bad}}$. Let the contaminated supply be $s_{\alpha}=\alpha s^{\prime}+(1-\alpha)s$.
@@
In Fig 2 we present an example of interference among $K=3$ submarkets. The market of interest is the perfectly separated market (right). This is because, in parallel A/B testing, submarkets are explicitly created such that each submarket resembles the global market. Then when a submarket receives a treatment, the observed quantities in that submarket, such as revenues and social welfare, are considered surrogates for the treatment effect in the global market. However, in practice we only observe the interfered finite market (left), which converges to the interfered limit market (middle). In App. F we show how to analyze parallel A/B testing using this framework.
4 A Debiased Estimator and Its Properties
@@
This section develops a methodology for making inferences about the uncontaminated limit FPPE. Since the interference setting is a special case of the contamination setting, we develop theories for the latter. We introduce a surrogate for pacing multipliers, based on the notion of directional derivatives, and establish its debiasing property in Sec. 4.1. Then, we focus on estimating this surrogate quantity in Sec. 4.2, and develop asymptotic normality results in Sec. 4.3. Secondly, we consider estimating revenue, which can be thought of as a smooth function of pacing multipliers. We discuss debiased revenue estimation and inference based on our pacing multiplier results in App. E.
4.1 A Debiased Surrogate for Pacing Multipliers